# Synthesize-on-Graph: Knowledgeable Synthetic Data Generation for Continue Pre-training of Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success but remain data-inefficient, especially when learning from small, specialized corpora with limited and proprietary data. Existing synthetic data generation methods for continue pre-training focus on intra-document content and overlook cross-document knowledge associations, limiting content diversity and depth. We propose Synthetic-on-Graph (SoG), a synthetic data generation framework that incorporates cross-document knowledge associations for efficient corpus expansion. SoG constructs a context graph by extracting entities and concepts from the original corpus, representing cross-document associations, and employing a graph walk strategy for knowledge-associated sampling. This enhances synthetic data diversity and coherence, enabling models to learn complex knowledge structures and handle rare knowledge. To further improve synthetic data quality, we integrate Chain-of-Thought (CoT) and Contrastive Clarifying (CC) synthetic, enhancing reasoning processes and discriminative power. Experiments show that SoG outperforms the state-of-the-art (SOTA) method in a multi-hop document Q&A dataset while performing comparably to the SOTA method on the reading comprehension task datasets, which also underscores the better generalization capability of SoG. Our work advances synthetic data generation and provides practical solutions for efficient knowledge acquisition in LLMs, especially in domains with limited data availability.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have achieved groundbreaking advancements in the field of Natural Language Processing (NLP), demonstrating the ability to acquire knowledge from unstructured text and perform complex, knowledge-intensive tasks Brown et al. (2020). These models have exhibited exceptional performance across various applications, including question-answering systems, machine translation, and conversational agents. This success is largely attributed to the next-word prediction objective Shannon (1951) combined with vast amounts of internet data Foundation (2007). However, despite these achievements, there remains a significant inefficiency in data utilization: while a 13-year-old human can acquire substantial knowledge from less than 100 million tokens Warstadt (2023), state-of-the-art open-source language models require training on 15 trillion tokens Dubey (2024).
This data inefficiency becomes particularly pronounced when models need to learn from small-scale, high-value corpora. With the increasing demand for proprietary domain knowledge, models are required to efficiently acquire information from limited data sources. For instance, in specialized fields such as medicine, law, or specific technological domains, the available data is not only limited but often proprietary. In such cases, traditional large-scale pretraining methods are inapplicable due to the unavailability of sufficient training data Kandpal (2023).
Moreover, recent studies have revealed limitations in the current pretraining paradigm. For example, models struggle when learning simple relations and require a large number of repeated instances to effectively learn facts Allen-Zhu & Li (2024). These issues become more acute when dealing with long-tail data or rare knowledge, as such information appears with extremely low frequency in large-scale corpora Kandpal (2023).
@@
To address the challenge of efficiently acquiring knowledge from small-scale corpora, synthetic data generation methods have been proposed for continue pretraining of models. They aim to expand the original limited data by generating diverse synthetic corpora, thereby improving the learning efficiency and performance of the models. For instance, the EntiGraph method decomposes the text corpus into a list of entities and generates descriptions about the relationships between entities, attempting to populate the underlying knowledge graph of the corpus Yang et al. (2024). However, as shown in Figure 1b this approach primarily focuses on intra-document content, neglecting inter-document knowledge associations. This leads to limitations in the content diversity and knowledge depth of the synthetic data. In reality, knowledge is often interconnected across documents and domains. Relying solely on entity combinations within a single document fails to capture the full spectrum of knowledge. Additionally, the lack of cross-document synthetic data constrains the model’s ability to handle complex, multi-hop problems that require integrating information from multiple documents to derive an answer. For instance, in the context graph in Figure 1a, the first encountered literature primarily describes Company X’s positive financial report and active market plans in 2023. However, relying on the across-document information associated with the entity "accounts receivable" —"companies with accounts receivable growth exceeding 30% face a special risk of bad debt" —we are able to derive a broader understanding of the literature: despite the positive net cash flow, people are suggested to be particularly cautious about the potential bad debt risk associated with Company X’s 35% accounts receivable growth. Cross-document information can integrate a multi-dimensional perspective on the topic (both positive and negative), build a progressive information chain, and uncover implicit phenomena, which realizes knowledge discovery where "1+1>2".
@@
To this end, we propose the Synthetic-on-Graph (SoG) framework—a context-graph-enhanced synthetic data generation method designed to provide an efficient solution for continue pretraining of LLMs. The core idea of SoG is to incorporate cross-document knowledge associations by constructing and leveraging a context graph to expand the original corpus effectively.
@@
Specifically, SoG comprises two key components: (1) Context Graph Construction and Cross-Document Sampling: We build a context graph from entities and concepts extracted from the original corpus, representing cross-document knowledge associations. Using this graph, we apply a two-stage cross-document sampling strategy: First random walk guided by document retrieval to achieve cross-document sampling, enhancing data diversity while preserving coherence and knowledge associations. This helps the model learn complex knowledge structures, especially for long-tail entities. Second, Secondary Sampling and Controlled Allocation help balance the knowledge distribution and support flexible data customization. (2)Combined Chain-of-Thought and Contrastive Clarifying Synthetic: We combine Chain-of-Thought and Contrastive Clarifying to enhance synthetic data quality. CoT guides the model to generate logical chains, improving depth and interpretability, while contrastive generation boosts the discriminative knowledge in the synthetic data.
@@
Through extensive experiments, we surpasses existing state-of-the-art (SOTA) methods on multi-hop document question answering in terms of effectiveness and achieve comparable performance on long-context reading comprehension datasets. Also, we demonstrate the better generalization capability over the SOTA method. The introduction of the SoG framework marks a significant advancement in synthetic data generation and continue pertaining (CPT) for LLMs, providing new directions and possibilities for future research. Our work not only drives the development of synthetic data techniques but also offers new perspectives for optimizing the training of LLMs.
@@ 2 Related Work
This section presents an overview of recent developments in synthetic data generation for the pretraining of large language models (LLMs). Synthetic data generation has emerged as a crucial area of research, with various strategies proposed to enhance the diversity and effectiveness of training datasets. A significant trend in this domain is the adoption of hierarchical prompting to generate targeted synthetic content. For instance, Eldan & Li (2023) utilize API-based LLMs to create children’s stories driven by specific keywords, illustrating that even smaller language models can yield fluent narratives when pre-trained on such datasets. Ma et al. (2024) achieve automatic analysis and annotation on complex data in the legal domain by using a modular multi-process pipeline, along with the injection of expert knowledge in the form of few-shot learning into each submodule. This approach was used for both pretraining and fine-tuning. This underscores the potential of hierarchical prompting in producing effective and relevant training data.
In another vein, Gunasekar et al. (2023) generate diverse educational content, such as textbooks and coding exercises, by conditioning on attributes like topic, audience, and function names. The datasets generated from this method have supported the development of robust LLMs, as further explored in subsequent studies (Li et al., 2023; Javaheripi et al., 2023). However, these approaches are often hindered by a lack of public accessibility to the datasets and prompt strategies, limiting reproducibility and broader community progress. Similarly, Maini et al. (2024) focus on rephrasing existing documents to generate new training data, reporting enhancements in training efficiency through these modified versions.
@@
Current efforts (Mecklenburg et al., 2024; Yang et al., 2025) explore synthetic QA generation for task-specific finetuning, reflect an emerging interest in incorporating knowledge-aware strategies into data generation. Although such strategies have demonstrated benefits for specific QA tasks, their applicability remains limited for more general-purpose tasks, indicating a gap that could potentially be filled by new data generation approaches that are untethered to any particular downstream application.
Moreover, Ovadia et al. (2023) explore continue pretraining of Llama 2 models using synthetic paraphrases of Wikipedia articles, with mixed results regarding performance improvements. This suggests limitations in relying solely on paraphrasing techniques to enhance model knowledge and underscores the need for research into more robust methods that can generate synthetic data with greater diversity and depth.
@@ 3.1.3 Context Graph
The context graph $\mathcal{G}$ is defined as $\mathcal{G}=(\mathcal{E},E)$, where: $\mathcal{E}$ represents the set of nodes corresponding to all entities, and
$E=\{(e_{x},e_{y})\mid\exists i,j\text{ such that }e_{x},e_{y}\in\mathcal{E}_{i,j}\}$
represents the set of edges signifying implicit contextual connections between entities.
@@ 3.2.1 Initialization
[…]
@@ 3.2.3 Secondary Sampling and Controlled Allocation
To further refine the control over synthetic data size, we iteratively allocate the secondary sampled paths into subsets according to the coverage of the original corpus, where each subset functions as an independent unit opted for the maximum corpus coverage ($>r$) and the most balanced paragraph frequency. This modular approach allows for seamless flexibility in data customization during the generation process: depending on the required volume of synthetic data, we can combine an appropriate number of subsets to support various experimental configurations. Specifically, due to the decreasing availability of sparse entities and associated texts as sampling iterations progress, the subset obtained in the first iteration should have the highest coverage $r$ of the original corpus. As the number of iterations increases, the coverage of subsequent subsets will gradually decrease under the fixed sampling size. We use the sample size of the first subset and corpus coverage $r$ as references and, based on the difference between the current iteration’s sampling rate $\triangle r=\frac{r-r^{\prime}}{r}$, re-sample and re-use texts of entities with the lowest utilization rate to complete the current path subset.
@@ 4 Experiments
To comprehensively evaluate the effectiveness and applicability of the proposed Synthetic-on-Graph (SoG) framework, this section explores its performance through a series of carefully designed experiments. The experiments aim to assess SoG’s contributions in four major aspects: First, to what extent does incorporating cross-document knowledge associations in SoG enhance the diversity and depth of synthetic data compared to intra-document-focused methods (RQ1)? Second, does SoG’s synthetic data provide consistent performance gains across language models of different sizes (RQ2)? Third, to what extent can SoG mitigate the long-tail knowledge problem in the original corpus (RQ3)? In what scenarios is SoG synthesis applicable? (RQ4)?
@@ 4.1 Datasets
To answer the RQs, we selected two datasets, MultiHop-RAG and a financial report corpus:
  • MultiHop-RAG (MHRAGTang & Yang (2024) is specifically designed to challenge the multi-hop reasoning capabilities of LLMs. It consists of queries constructed from news articles published between September and December 2023, which include information beyond the training cutoff of existing LLMs, ensuring that synthetic data is required to fill knowledge gaps. In addition, each query requires models to integrate evidence from multiple documents, mimicking real-world scenarios where knowledge is dispersed across sources. Existing LLMs, even RAG systems, often struggle with such tasks, underperforming in tasks that demand integrating and reasoning over scattered evidence. This dataset serves as an ideal benchmark to evaluate how SoG-generated synthetic data equips LLMs to utilize their internal knowledge for handling complex multi-hop reasoning effectively.
  • QuALITY Pang et al. (2022) is a multiple-choice question-answering dataset for long document comprehension. Unlike in prior work with passages, the questions are written and validated by contributors who have read the entire passage, rather than relying on summaries or excerpts. For a fair comparison with the state-of-the-art CPT synthetic data method, EntiGraph, we also chose this dataset for evaluation.
@@ 4.2 Baselines and Metrics
We choose Direct QA (LLM zero-shot answering) and the state-of-the-art methods EntiGraph Yang et al. (2024) as baselines for evaluation. The evaluation metrics for MHRAG and QuALITY are Exact Match (EM) and Accuracy (Acc) respectively.
@@ 4.3 Experiment Details
In our generation setup, we used GPT-4o-mini as the generation model. The temperature was set at 0.7. We utilize semantic chunking [^1] https://python.langchain.com/docs/how_to/semantic-chunker/ to split the long contexts. The semantic embedding was computed by bge-small-en-v1.5. In all experiments, we continue pretrain the LLMs with a context length of 2048 and a batch size of 64. We apply a linear learning rate warmup for 10% of the total steps, followed by a cosine decay with a peak learning rate of 5e-6. We perform full-parameter training for 2 epochs in BF16 precision, using a per-device batch size of 2 and accumulating gradients over 4 steps. In addition, within 4.5×4.5$\times$4.5 × of the original corpus size, the sampling paths for CoT are of one-hop length, while beyond that, the sampling paths are of two-hop length. For QuALITY, we followed the evaluation setup in EntiGraph. For MHRAG, we evaluate the CPT models with zero-shot prompting on a sample of 1,000 QA pairs.
@@ 4.4 Main Experiment Results
To answer RQ1 and RQ2, we compare the effectiveness of SoG and the intra-document-focused method EntiGraph in continue pre-training (CPT) with varying amounts of synthetic data on two datasets. The results are shown in Figure 3. For MHRAG, model performance steadily improves as the amount of SoG synthetic data increases. In contrast, EntiGraph synthetic data provides limited gains. When the EntiGraph data size exceeds 1.5 times the original corpus, performance plateaus or even degrades due to its reliance on intra-document associations. This limitation prevents diverse and deeper generations, especially for complex tasks requiring cross-source knowledge integration. The sharp performance gap on MHRAG underscores the strength of SoG’s cross-document knowledge integration in the context graph, which uncovers implicit entity relationships and enables richer reasoning. In addition, the most significant performance boost from SoG occurs when the synthetic data volume is within 0 to 1.5 times the original corpus, demonstrating that even a moderate amount of SoG data effectively enhances large model performance.
@@
00.71.534.569234050607080Synthetic Size (Multiples of Original Corpus)Score (%)SoG (MHRAG)EntiGraph (MHRAG)SoG (QuALITY)EntiGraph (QuALITY)Raw CPT (MHRAG)Raw CPT (QuALITY)
Figure 3: Performance trends of SoG and EntiGraph.

### v3

# Synthesize-on-Graph: Knowledgeable Synthetic Data Generation for Continued Pre-training of Large Language Models

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success but remain data-inefficient, especially when learning from small, specialized corpora with limited and proprietary data. Existing synthetic data generation methods for continue pre-training focus on intra-document content and overlook cross-document knowledge associations, limiting content diversity and depth. We propose Synthetic-on-Graph (SoG), a synthetic data generation framework that incorporates cross-document knowledge associations for efficient corpus expansion. SoG constructs a context graph by extracting entities and concepts from the original corpus, representing cross-document associations, and employing a graph walk strategy for knowledge-associated sampling. This enhances synthetic data diversity and coherence, enabling models to learn complex knowledge structures and handle rare knowledge. To further improve the quality of synthetic data, we integrate two complementary strategies, Chain-of-Thought (CoT) and Contrastive Clarifying (CC), to enhance both reasoning capability and discriminative power. Extensive experiments demonstrate that SoG surpasses state-of-the-art (SOTA) methods on multi-hop and domain-specific question answering, while achieving competitive performance on long-context reading comprehension. These results highlight the superior generalization ability of SoG. Our work advances the paradigm of synthetic data generation and offers practical solutions for efficient knowledge acquisition in LLMs, particularly for downstream tasks and domains with limited training data.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have achieved groundbreaking advancements in the field of Natural Language Processing (NLP), demonstrating the ability to acquire knowledge from unstructured text and perform complex, knowledge-intensive tasks [1]. These models have exhibited exceptional performance across various applications, including question-answering systems, machine translation, and conversational agents. This success is largely attributed to the next-word prediction objective [2] combined with vast amounts of internet data [3]. However, despite these achievements, there remains a significant inefficiency in data utilization [4].
This data inefficiency becomes particularly pronounced when models need to learn from small-scale, high-value corpora. With the increasing demand for proprietary domain knowledge, models are required to efficiently acquire information from limited data sources. For instance, in specialized fields such as medicine, law, or specific technological domains, the available data is not only limited but often proprietary. In such cases, traditional large-scale pretraining methods are inapplicable due to the unavailability of sufficient training data [5].
Moreover, recent studies have revealed limitations in the current pretraining paradigm. For example, models struggle when learning simple relations and require a large number of repeated instances to effectively learn facts [6]. These issues become more acute when dealing with long-tail data or rare knowledge, as such information appears with extremely low frequency in large-scale corpora [5].
@@
To address the challenge of efficiently acquiring knowledge from small-scale corpora, synthetic data generation methods have been proposed for continued pretraining of models. They aim to expand the original limited data by generating diverse synthetic corpora, thereby improving the learning efficiency and performance of the models. For instance, the EntiGraph method decomposes the text corpus into a list of entities and generates descriptions about the relationships between entities, attempting to populate the underlying knowledge graph of the corpus [7]. However, as shown in Figure 1b this approach primarily focuses on intra-document content, neglecting inter-document knowledge associations. This leads to limitations in the content diversity and knowledge depth of the synthetic data. In reality, knowledge is often interconnected across documents and domains. Relying solely on entity combinations within a single document fails to capture the full spectrum of knowledge. Additionally, the lack of cross-document synthetic data constrains the model’s ability to handle complex, multi-hop problems that require integrating information from multiple documents to derive an answer. For instance, in the context graph in Figure 1a, the first encountered literature primarily describes Company X’s positive financial report and active market plans in 2023. However, relying on the across-document information associated with the entity "accounts receivable" —"companies with accounts receivable growth exceeding 30% face a special risk of bad debt" —we can derive a broader understanding of the literature: despite the positive net cash flow, people are suggested to be particularly cautious about the potential bad debt risk associated with Company X’s 35% accounts receivable growth. Cross-document information can integrate multi-dimensional perspectives on a topic (both positive and negative), build a progressive chain of information, and uncover implicit phenomena — integrating knowledge in a way that uncovers more than what each document alone can offer, where "1+1>2".
@@
To this end, we propose the Synthesize-on-Graph (SoG) framework—a context-graph-enhanced synthetic data generation method designed to provide an efficient solution for continued pretraining of LLMs. The core idea of SoG is to incorporate cross-document knowledge associations by constructing and leveraging a context graph to expand the original corpus effectively.
@@
Specifically, SoG comprises two key components: (1) Context Graph Construction and Cross-Document Sampling: We build a context graph from entities and concepts extracted from the original corpus, representing cross-document knowledge associations. Using this graph, we apply a two-stage cross-document sampling strategy: first, random walks guided by document retrieval to achieve cross-document exploration, enhancing data diversity while preserving coherence and knowledge associations. This helps the model learn complex knowledge structures, especially for long-tail entities. Second, Secondary Sampling and Controlled Allocation help balance the knowledge distribution and support flexible data customization. (2)Combined Chain-of-Thought and Contrastive Clarifying Synthesize: We combine Chain-of-Thought and Contrastive Clarifying to enhance synthetic data quality. CoT guides the model to generate logical chains, improving depth and interpretability, while contrastive generation boosts the discriminative knowledge in the synthetic data.
@@
Through extensive experiments, our approach outperforms existing state-of-the-art (SOTA) methods on multi-hop document and domain-specific question answering tasks, while achieving comparable results on long-context reading comprehension. Also, we demonstrate better generalization capability over the SOTA method. The introduction of the SoG framework marks a significant advancement in synthetic data generation and continued pertaining (CPT) for LLMs, providing new directions and possibilities for future research. Our work not only drives the development of synthetic data techniques but also offers new perspectives for optimizing the training of LLMs.
@@ 2 Related Work
This section presents an overview of recent developments in synthetic data generation for the pretraining of large language models (LLMs). Synthetic data generation has emerged as a crucial area of research, with various strategies proposed to enhance the diversity and effectiveness of training datasets. A significant trend in this domain is the adoption of hierarchical prompting to generate targeted synthetic content. For instance, [8] utilize API-based LLMs to create children’s stories driven by specific keywords, illustrating that even smaller language models can yield fluent narratives when pre-trained on such datasets. [9] achieve automatic analysis and annotation on complex data in the legal domain by using a modular multi-process pipeline, along with the injection of expert knowledge in the form of few-shot learning into each submodule. This approach was used for both pretraining and fine-tuning. This underscores the potential of hierarchical prompting in producing effective and relevant training data.
In another vein, [10] generate diverse educational content, such as textbooks and coding exercises, by conditioning on attributes like topic, audience, and function names. The datasets generated from this method have supported the development of robust LLMs, as further explored in subsequent studies [11, 12]. However, these approaches are often hindered by a lack of public accessibility to the datasets and prompt strategies, limiting reproducibility and broader community progress. Similarly, [13] focus on rephrasing existing documents to generate new training data, reporting enhancements in training efficiency through these modified versions.
@@
Current efforts [14, 15] explore synthetic QA generation for task-specific finetuning, reflecting an emerging interest in incorporating knowledge-aware strategies into data generation. Although such strategies have demonstrated benefits for specific QA tasks, their applicability remains limited for more general-purpose tasks, indicating a gap that could potentially be filled by new data generation approaches that are untethered to any particular downstream application.
Moreover, [16] explore continued pretraining of Llama 2 models using synthetic paraphrases of Wikipedia articles, with mixed results regarding performance improvements. This suggests limitations in relying solely on paraphrasing techniques to enhance model knowledge and underscores the need for research into more robust methods that can generate synthetic data with greater diversity and depth.
@@ 3 Methodology
$\text{Corpus}\xrightarrow{\text{Construction}}\text{Context Graph}\xrightarrow{\text{Context-graph Traversal}}\text{Path Set }\mathcal{P}$ (1)
$\xrightarrow{\text{Secondary Sampling}}\text{Balanced Path Set }\mathcal{P^{*}}\xrightarrow{\text{Generation Strategies}}\text{Synthetic Data}$
@@ 3.1.3 Context Graph
We define a context graph $\mathcal{G}=(\mathcal{E},E)$, where $\mathcal{E}$ denotes the set of nodes corresponding to all identified entities. The edge set is given by
$E=\{(e_{x},e_{y})\mid\exists i,j\;\;\text{s.t.}\;\;e_{x},e_{y}\in\mathcal{E}_{i,j}\},$
where $\mathcal{E}_{i,j}$ represents the subset of entities co-occurring within a bounded textual unit (e.g., a paragraph or sentence). Thus, an edge between $e_{x}$ and $e_{y}$ is induced whenever the two entities are observed to co-occur within the same discourse context. In this way, the graph topology captures implicit contextual associations among entities, with co-occurrence serving as a distributional proxy for semantic relatedness.
@@ 3.2.1 Initialization
To enhance content diversity and knowledge association across multiple documents, we implement a cross-document sampling strategy that traverses the constructed context graph $\mathcal{G}=(\mathcal{G},\mathcal{R})$. Starting from a root entity $e_{root}\in\mathcal{E}$, we perform a breadth-first search (BFS) traversal to collect multi-hop paths that link related entities and their associated text paragraphs across documents. We will traverse all nodes in $\mathcal{G}$ as a root entity.
@@ 3.2.3 Secondary Sampling and Controlled Allocation
To further refine the control over synthetic data size, we iteratively allocate the secondary sampled paths into subsets according to the coverage of the original corpus, where each subset functions as an independent unit opted for the maximum corpus coverage (>r>r) and the most balanced paragraph frequency. This modular approach allows for seamless flexibility in data customization during the generation process: depending on the required volume of synthetic data, we can combine an appropriate number of subsets to support various experimental configurations. Specifically, due to the decreasing availability of sparse entities and associated texts as sampling iterations progress, the subset obtained in the first iteration should have the highest coverage $r$ of the original corpus. As the number of iterations increases, the coverage of subsequent subsets will gradually decrease under the fixed sampling size. We use the sample size of the first subset and corpus coverage $r$ as references and, based on the difference between the current iteration’s sampling rate $\triangle r=\frac{r-r^{\prime}}{r}$, re-sample and re-use texts of entities with the lowest utilization rate to complete the current path subset.
@@ 4 Experiments
To comprehensively evaluate the effectiveness and applicability of the proposed Synthesize-on-Graph (SoG) framework, this section explores its performance through a series of carefully designed experiments. The experiments aim to assess SoG’s contributions in four major aspects: First, to what extent does incorporating cross-document knowledge associations in SoG enhance the diversity and depth of synthetic data compared to intra-document-focused methods (RQ1)? Second, does SoG’s synthetic data provide consistent performance gains across language models of different sizes (RQ2)? Third, to what extent can SoG mitigate the long-tail knowledge problem in the original corpus (RQ3)? In what scenarios is SoG synthesis applicable? (RQ4)?
@@ 4.1 Datasets
To address our research questions, we evaluate on three representative datasets: MultiHop-RAG, BioASQ, and QuALITY. A detailed description of each dataset is provided in the Appendix A.1.
@@ 4.2 Baselines and Metrics
We choose Direct QA (directly answering by the base model), Rephrasing (back-translation and synonym replacement, following [17]) and the state-of-the-art methods EntiGraph [7] as baselines for evaluation. The evaluation metrics for MHRAG, BioASQ and QuALITY are Exact Match (EM), model-based evaluation (MBE) approach using LLM-as-a-Judge[18], and Accuracy (Acc), respectively.
@@ 4.3 Experiment Details
In our generation setup, we used GPT-4o-mini as the generation model. The temperature was set at 0.7. We utilize semantic chunking [^1] https://python.langchain.com/docs/how_to/semantic-chunker/ to split the long contexts. The semantic embedding was computed by bge-small-en-v1.5. In all experiments, we continued pretrain the LLMs with a context length of 2048 and a batch size of 64. We apply a linear learning rate warmup for 10% of the total steps, followed by a cosine decay with a peak learning rate of 5e-6. We perform full-parameter training for 2 epochs in BF16 precision, using a per-device batch size of 2 and accumulating gradients over 4 steps. In addition, within $4.5\times$ of the original corpus size, the sampling paths for CoT are of one-hop length, while beyond that, the sampling paths are of two-hop length. For QuALITY, we followed the evaluation setup in EntiGraph. For MHRAG, we evaluate the CPT models with zero-shot prompting on a sample of 1,000 QA pairs. For BioASQ, we constructed a hard subset consisting of 1,114 questions that Qwen3-8B failed to answer correctly in a single attempt. This sampling criterion ensures that the selected questions reflect genuine challenges for strong LLMs, thereby providing a more rigorous evaluation of knowledge-intensive reasoning. For entity ambiguity issue, we rely on surface-form string matching combined with simple heuristics, including normalization of singular/plural forms and letter casing, alias matching, and Wikipedia-style redirect mappings to partially address this issue.
@@ 4.4 Main Experiment Results
To answer RQ1 and RQ2, we compare the effectiveness of SoG, traditional Rephrasing augmentation and the intra-document-focused method EntiGraph in continued pre-training (CPT) with varying amounts of synthetic data on two datasets. The results are shown in Figure 3. For MHRAG and BioASQ, model performance steadily improves as the amount of SoG synthetic data increases. In contrast, EntiGraph synthetic data provides limited gains. Especially in MHRAG, when the EntiGraph data size exceeds 1.5 times the original corpus, performance plateaus or even degrades due to its reliance on intra-document associations. This limitation prevents diverse and deeper generations, especially for complex tasks requiring cross-source knowledge integration. The sharp performance gap on MHRAG underscores the strength of SoG’s cross-document knowledge integration in the context graph, which uncovers implicit entity relationships and enables richer reasoning. In addition, the most significant performance boost from SoG occurs when the synthetic data volume is within 0 to 1.5 times the original corpus, demonstrating that even a moderate amount of SoG data effectively enhances large model performance.
@@
Moreover, the traditional Rephrasing augmentation method yields only marginal or negligible improvements across all datasets, further highlighting the necessity of structurally informed synthetic data construction.
Finally, another observation is that CPT solely on the original corpus yields at best limited gains and in some cases even degrades performance relative to the original model (see Raw CPT in Figure 3). We attribute this to the lack of diversity and distributional differences in the original corpus, which further emphasizes the critical role of Synthetic CPT.
00.71.534.5692310102020303040405050606070708080Synthetic Size (Multiples of Original Corpus)Score (%)SoG (MHRAG)SoG (QuALITY)SoG (BioASQ)EntiGraph (MHRAG)EntiGraph (QuALITY)EntiGraph (BioASQ)Rephrasing (MHRAG)Rephrasing (QuALITY)Rephrasing (BioASQ)Raw CPT (MHRAG)Raw CPT (QuALITY)Raw CPT (BioASQ)
Figure 3: Performance trends of SoG and EntiGraph across three benchmarks with rephrasing and CPT baselines.